# Hot Jupiters are Inflated Primarily by Shallow Heating

### v1

@@ 3 Host Star & Planet Property Inference
    Galaxy Evolution Explorer (GALEX; mart05) near-ultraviolet (NUV) photometry from the GUVcat_AIS (bia17), or, if the system is in the Kepler field, the GALEX-CAUSE Kepler (GCK) survey (olm15) including in quadrature a zero-point uncertainty of 0.02 mag in both cases;
@@
    SkyMapper Southern Survey Data Release (DR) 4 uvgrizuvgriz photometry including in quadrature their zero-point uncertainties (0.03, 0.02, 0.01, 0.01, 0.01, 0.02) mag (onk24);
@@
    the most up-to-date Sloan Digital Sky Survey (SDSS) ugrizugriz photometry from DR13 including in quadrature their zero-point uncertainties (0.015, 0.003, 0.003, 0.003, 0.003) mag (fuk96; gun98; gun06; yor00; doi10; fin16; alb17);
@@
    Panoramic Survey Telescope & Rapid Response System (Pan-STARRS) DR2 grizygrizy photometry including in quadrature their zero-point uncertainties (0.014, 0.014, 0.015, 0.015, 0.018) mag (cha16; fle20; mag20a; mag20b; mag20c; wat20);
@@
    Kepler Input Catalog (KIC) grizgriz photometry (if the system is in the Kepler field) including in quadrature their zero-point uncertainties of about 0.03 mag (bro11);
@@
    Tycho-2 $B_{T}$ and $V_{T}$ photometry including in quadrature their zero-point uncertainties (0.078, 0.058) mag (hog00; marz05);
@@
    Gaia EDR3 $G$ photometry including in quadrature its zero-point uncertainty (Gaia_mission2016; GaiaEDR3; GaiaEDR3Validation; EDR3photometry; row21; tor21);
@@
    Two-micron All-sky Survey (2MASS) $JHK_{s}$ photometry including their zero-point uncertainties (skr06);
@@
    Wide-field Infrared Survey Explorer (WISE) All-sky (wri10) or AllWISE (mai11) $W1$ and $W2$ photometry in quadrature their zero-point uncertainties (0.032,0.037) mag [^3] https://wise2.ipac.caltech.edu/docs/release/allsky/expsup/sec4_4h.html#PhotometricZP. Because the procedures used in the construction of the WISE All-sky catalog were better for saturated sources than the procedures used in the construction of the AllWISE catalog, we use WISE All-sky data when $2\leq W1<8$ and $1.5\leq W2<7$. We use AllWISE data when $W1\geq 8$ and $W2\geq 7$.
@@
    We also fit to a zero point-corrected Gaia EDR3 parallax (GaiaEDR3; GaiaEDR3Validation; lin21a; lin21b; row21; tor21) and
@@
    when available an estimated extinction value based on the gre19 three-dimensional extinction map calculated using the dustmaps Python module (gre18). If an extinction estimate is unavailable from that source, then we use a different three-dimensional extinction map (lal22; ver22).
We use the data quality flags described in Hamer22. We only include ground-based optical data from one survey in each fit, prioritizing the survey that provides ultraviolet data. If more than one survey provides ultraviolet data, then use data from the survey that provides the most independent flux measurements. If no survey provides ultraviolet data, then we use data from the survey that provides the most independent flux measurements.
For our priors, we use a cha03 log-normal mass prior for $M_{\ast}<1~M_{\odot}$ joined to a sal55 power-law prior for $M_{\ast}\geq 1~M_{\odot}$, a metallicity prior based on the Geneva-Copenhagen Survey (GCS; cas11), a log-uniform age prior between 100 Myr and the age of the universe 13.721 Gyr, a uniform extinction prior in the interval between the value described above minus/plus five times its uncertainty, and a distance prior proportional to volume between the bai21 geometric distance minus/plus five times its uncertainty.
Next, we obtain planet parameters for the exoplanets in systems for which an isochrones stellar property inference was successful. We select our sample of exoplanet system measurements from the NASA Exoplanet Archive’s ps table (ExoplanetArchive; Christiansen25; PS). We eliminate systems discovered by pulsar timing, microlensing, and direct imaging as planet radii cannot be inferred for these systems. Our process for calculating planet parameters prefers to use as few Exoplanet Archive entries as possible for each planet’s orbit solution (Doppler semiamplitude $K$ and eccentricity $e$) and transit solution (orbital period $P_{\text{orb}}$, inclination $i$ or impact parameter $b$, and planet-to-star radius ratio $q$ or transit depth $\delta$). We prefer transit solutions with inclinations over impact parameters and radius ratios over transit depths as inclinations and radius ratios do not require additional transformations to obtain a mass or radius inference. In cases where a transit depth is reported instead of a planet-to-star radius ratio, we transform the transit depth radius ratio by taking its square root and transform its uncertainty in quadrature. We choose the larger of the upper and lower uncertainty for each parameter in cases where they are unequal as the inputs to our hierarchical Bayesian inference must be normally distributed.
@@
In a similar fashion to Schmidt23, we re-parameterize Equation (6) as a cubic equation
@@
4 Hierarchical Bayesian Inference
@@
$^{i}, 𝐡\mathbf{h}$ Heat factor (see Section 4)
@@
Note. — The superscript $i$ indicates the $i^{\mathrm{th}}$ planet in our sample, and bolded variables are the vector of that parameter for all planets. The covariance $\mathbf{\Sigma_{*}^{i}}$ is output by the isochrones run for each host star. Additionally, as the MIST grids used by isochrones lack the initial bulk metallicity column, we fit for this using the stellar age, mass, and present-day photospheric metallicity to obtain the initial bulk metallicity for each system (see Section 4.1 for more details).
To handle possible variations in the heating efficiency $\epsilon(F)$ (Eq. 3), we parameterize the variation with $h^{i}$, which is the offset relative to the mean heating in number of standard deviations. To match the results of Thorngren2018 exactly, we could set the prior as $h\sim\mathcal{N}(0,1^{2})$ where $\mathcal{N}(\mu,\sigma^{2})$ is a normal distribution with mean $\mu$ and standard deviation $\sigma$. However, to allow the model additional flexibility while maintaining the functional form of $\epsilon(F)$ we instead allow a global mean offset to the heating $H$, such that $h^{i}\sim\mathcal{N}(H,1^{2})$. Due to some differences in equation of state between this article and Thorngren2018, we expect $H$ to be somewhat negative (i.e. less heating is required).
@@ 4.1 Sampling Strategy
Note. — This table is sorted by planet name and is available in machine-readable format. The sample enumerated by this table has not been vetted for population-level biases and should not be used for occurrence analyses.
Our thermal evolution model’s stellar model grid is adapted from the MIST grid and takes the star’s initial bulk metallicity as an input. However, the MIST model grid used by the isochrones Python package is restricted to a subset of parameters that does not include it; as a result the posteriors instead report the present-day photospheric metallicity. This requires us to fit the initial bulk metallicity posterior for each star from its inferred present-day photospheric metallicity, stellar mass, and stellar age posteriors. While this process works with no issue for the vast majority of the host stars in our sample, some metal-rich stars lack suitable MIST models at a high enough initial bulk metallicity to match the inferred present-day photospheric metallicity from their isochrones posterior. Performing the initial bulk metallicity inference procedure cuts off a significant portion of these stars’ metallicity posteriors. We therefore exclude any stars for which this occurs to more than a few percent of samples.
We perform individual interior structure retrievals on all of the planets in our sample for which our grid of models is suitable: 0.1 $_{\text{Jup}}<M^{i}_{\text{obs}}<4~M$$_{\text{Jup}}$, 0.5 $_{\odot}<M^{i}_{\ast\text{obs}}<2~M$, and $0.75<\text{[Fe/H]}^{i}_{\text{0}}<0.5$. These resulting posterior distributions are often non-Gaussian, so to facilitate convergence we utilize a two-step MCMC process. We first run a shorter 20,000-iteration initial retrieval with a 5000-iteration burn-in to characterize the covariance, and then run a second 250,000-iteration retrieval with a 5000-iteration burn-in and using the covariance of the initial run to set the proposal covariance. We apply a thinning factor of 10 to both of these steps, where we only record a sample every 10 iterations. In some cases the model fails to converge or otherwise fails to provide plausible results; we exclude these planets from the upper level of the hierarchical model. We list the planets in our sample for which we successfully retrieve the interior structure following this methodology in Table 2.
@@
5 Results and Discussion
We show an example corner plot from retrievals on WASP-6 b (green) and CoRoT-25 b (blue) in Figure 3. Our model allows a relatively young planet such as WASP-6 b to be larger due to either delayed cooling or a less dense envelope caused by a lower bulk metallicity. This manifests in the posterior as a degeneracy between the planet’s bulk metallicity $Z_{p}$ and intrinsic luminosity reduction factor $f_{\text{L}}$. On the other hand, a planet like CoRoT-25 b (Fig. 3, right) is too old to be affected by delayed cooling, having already reached its equilibrium size (see Sec. 2). Such planets are immune to the $Z_{p}$-$f_{\text{L}}$ degeneracy, making them a natural control group for delayed cooling and useful for constraining the population mean heat factor $H$. These examples show that a single planet cannot provide enough information to fully constrain $f_{\text{L}}$, making a population-level analysis necessary.
@@ 5 Results and Discussion
Next, we show the results of our hierarchical Bayesian inferences in Figure 4. Each corner plot in Figure 4 shows the posterior and covariance for our population mean heat factor $H$ and intrinsic luminosity reduction factor $f_{\text{L}}$. These parameters are uncorrelated, indicating that the delayed cooling effect we see is not biased by or an artifact of our thermal evolution model’s equation of state choice. We vary the value of each retrieved planet’s initial specific entropy from $s_{\text{init}}=11~k_{B}/\text{baryon}$ to $s_{\text{init}}=9~k_{B}/\text{baryon}$ in increments of $0.5~k_{B}/\text{baryon}$. These represent the range between “hot start” and “cold start” giant planet formation models (Marley2007) to avoid limiting our modeling to a particular assumption about how the planet initially formed. Nevertheless, our results are consistent with one another for each value of $s_{\text{init}}$.
@@
Figure 4: Corner plots for the hyperparameters of our population-level intrinsic luminosity reduction factor $f_{\text{L}}$ and population mean heat factor $H$. Each posterior’s model assumes a different initial specific entropy for the individual interior structure retrievals it marginalizes over. Not shown are the individual planets’ heat factor hyperparameters, which are included to account for small differences in the theoretical equation of state used between this work and Thorngren2018. We find that 2-3% of the heating hot Jupiters experience can be attributed to deep heating, regardless of our assumption for the initial specific entropy.
@@
We find that the interior cooling rates of the hot Jupiter population are reduced by a factor of between 20 and 60 relative to the simpler heating model presented in Thorngren2018. In cases where the planet has reached thermal equilibrium, this is equivalent to shallow heating accounting for about 97% of the total heating experienced by hot Jupiters [^4] Note that while most of the planets in our sample are not exactly in equilibrium, the older ones are not far off from it.. In other words, the delayed cooling we observe in the hot Jupiter population means that the youngest ($t^{i}<$1 Gyr) planets in our sample are consistently larger than the trend inferred by Thorngren2018. We additionally find that the population mean heat $H$ (see Sec. 4) is modest but nonzero to correct for the updated equation of state. This means that the base amount of heating required to inflate hot Jupiters’ radii is slightly lower than Thorngren2018. The lack of correlation between $H$ and $f_{\text{L}}$ in our posteriors (Fig. 4) shows that this was not a bias on the latter.
@@
The high intrinsic luminosity reduction factor we find is best explained by predominantly shallow heating (near the surface, rather than the deep interior). The proposed hot Jupiter heating mechanisms that would cause shallow heating, including Ohmic dissipation (Batygin2010; Batygin2011), turbulent mixing (Youdin2010), and longitudinal advection (Tremblin2017) preferentially occur near radiative-convective boundary. While a larger-than-expected opacity in the shallow interior (Burrows2003) may also contribute, the extent to which our hot Jupiters’ radii are inflated at the population level indicates that this explanation alone is insufficient. Similarly, strong compositional gradients (e.g., Chabrier2007; Leconte2012) would be insufficient, as neither they nor enhanced opacities could explain the flux-heating relation of Thorngren2018. As a result, while we cannot definitively rule out these alternatives entirely, they can only be minor contributions compared to shallow heating mechanisms.
@@
Our results strongly exclude $f_{\text{L}}=1$, meaning some delay is required to explain the larger radii of young hot Jupiters. On the other hand, exclusively delayed cooling effects from shallow heating would not reinflate planets as seen in hot Jupiters orbiting main sequence (Hartman2016) and evolved (Grunblatt2016; Grunblatt2017) stars. The values of $f_{\text{L}}$ our model prefers, a few percent of standard cooling rates, instead balance the need for both inflated young hot Jupiters and reinflation of older ones. This resolves the tension between the deep heating models of Thorngren2018 and the delayed cooling seen in Thorngren2021.
@@
Dominant shallow heating has implications for the relationship between observed phase curves and radius inflation. This is because the strength of many shallow heating mechanisms is coupled to wind speeds in the atmosphere (see Menou2012), so circulation-related observables should mirror the heating efficiency-equilibrium temperature trend found in Thorngren2018. Specifically, we expect hot spot offsets for hot Jupiters to scale with the log of the incident flux as $\exp{\left[(\log(F)-\log(F_{0}))^{2}/(2w)\right]}$, where $F_{0}\approx 1.38$ Gerg s-1 cm-2 and $w\approx 0.37$ (Thorngren2024). This corresponds to a peak hot spot offset at $T_{\mathrm{eq}}\approx 1600$ K, decreasing again at higher temperatures. Though some population studies of archival Spitzer phase curves exist (e.g., May22; Dang25), they are limited in sample size as well as the range of equilibrium temperatures and instellations they probe. Future observing campaigns, as well as missions like European Space Agency’s Ariel mission (Tinetti22), could observe a larger statistical sample to evaluate these predictions.
@@
While our sample was large enough to obtain definitive results, a larger sample of younger ($t^{i}_{\text{obs}}\lesssim 1$ Gyr) planets with precise masses, radii, and host star properties would enable testing of more complex thermal evolution models. On the other end of the exoplanet age distribution, a larger sample of hot Jupiters orbiting evolved stars could help constrain the timescale for reinflation and therefore the effects of deep heating thanks to their much faster rate of stellar evolution. The careful study of higher-order effects due to tidal circularization and inspiral (e.g., Jackson2008; Leconte2010) and/or late-arrival scenarios (e.g., Hamer22, ; Schmidt et al. 2025 submitted) further altering the planet’s interior heating or instellation could also provide more physically-motivated constraints on delayed cooling. These larger and more precise data sets would enable further characterization of the mechanisms that heat hot Jupiters, leading to much more informed constraints on individual mechanism contributions as well as their typical impacts on observed planetary atmosphere characteristics.
6 Conclusions

### v2

@@ III Host Star & Planet Property Inference
    Galaxy Evolution Explorer (GALEX; Martin et al., 2005) near-ultraviolet (NUV) photometry from the GUVcat_AIS (Bianchi et al., 2017), or, if the system is in the Kepler field, the GALEX-CAUSE Kepler (GCK) survey (Olmedo et al., 2015) including in quadrature a zero-point uncertainty of 0.02 mag in both cases;
@@
    SkyMapper Southern Survey Data Release (DR) 4 uvgrizuvgriz photometry including in quadrature their zero-point uncertainties (0.03, 0.02, 0.01, 0.01, 0.01, 0.02) mag (Onken et al., 2024);
@@
    the most up-to-date Sloan Digital Sky Survey (SDSS) ugrizugriz photometry from DR13 including in quadrature their zero-point uncertainties (0.015, 0.003, 0.003, 0.003, 0.003) mag (Fukugita et al., 1996; Gunn et al., 1998, 2006; York et al., 2000; Doi et al., 2010; Finkbeiner et al., 2016; Albareti et al., 2017);
@@
    Panoramic Survey Telescope & Rapid Response System (Pan-STARRS) DR2 grizygrizy photometry including in quadrature their zero-point uncertainties (0.014, 0.014, 0.015, 0.015, 0.018) mag (Chambers et al., 2016; Flewelling et al., 2020; Magnier et al., 2020a, c, b; Waters et al., 2020);
@@
    Kepler Input Catalog (KIC) grizgriz photometry (if the system is in the Kepler field) including in quadrature their zero-point uncertainties of about 0.03 mag (Brown et al., 2011);
@@
    Tycho-2 $B_{T}$ and $V_{T}$ photometry including in quadrature their zero-point uncertainties (0.078, 0.058) mag (Høg et al., 2000; Maíz-Apellániz, 2005);
@@
    Gaia EDR3 $G$ photometry including in quadrature its zero-point uncertainty (Gaia Collaboration et al., 2016, 2021; Fabricius et al., 2021; Riello et al., 2021; Rowell et al., 2021; Torra et al., 2021);
@@
    Two-micron All-sky Survey (2MASS) $JHK_{s}$ photometry including their zero-point uncertainties (Skrutskie et al., 2006);
@@
    Wide-field Infrared Survey Explorer (WISE) All-sky (Wright et al., 2010) or AllWISE (Mainzer et al., 2011) $W1$ and $W2$ photometry in quadrature their zero-point uncertainties (0.032,0.037) mag [^5] https://wise2.ipac.caltech.edu/docs/release/allsky/expsup/sec4_4h.html#PhotometricZP. Because the procedures used in the construction of the WISE All-sky catalog were better for saturated sources than the procedures used in the construction of the AllWISE catalog, we use WISE All-sky data when $2\leq W1<8$ and $1.5\leq W2<7$. We use AllWISE data when $W1\geq 8$ and $W2\geq 7$.
@@
    We also fit to a zero point-corrected Gaia EDR3 parallax (Gaia Collaboration et al., 2021; Fabricius et al., 2021; Lindegren et al., 2021a, b; Rowell et al., 2021; Torra et al., 2021) and
@@
    when available an estimated extinction value based on the Green et al. (2019) three-dimensional extinction map calculated using the dustmaps Python module (Green, 2018). If an extinction estimate is unavailable from that source, then we use a different three-dimensional extinction map (Lallement et al., 2022; Vergely et al., 2022).
We use the data quality flags described in Hamer and Schlaufman (2022). We only include ground-based optical data from one survey in each fit, prioritizing the survey that provides ultraviolet data. If more than one survey provides ultraviolet data, we then use data from the survey that provides the most independent flux measurements. If no survey provides ultraviolet data, then we use data from the survey that provides the most independent flux measurements.
For our priors, we use a Chabrier (2003) log-normal mass prior for $M_{\ast}<1~M_{\odot}$ joined to a Salpeter (1955) power-law prior for $M_{\ast}\geq 1~M_{\odot}$, a metallicity prior based on the Geneva-Copenhagen Survey (GCS; Casagrande et al., 2011), a log-uniform age prior between 100 Myr and the age of the universe 13.721 Gyr, a uniform extinction prior in the interval between the value described above minus/plus five times its uncertainty, and a distance prior proportional to volume between the Bailer-Jones et al. (2021) geometric distance minus/plus five times its uncertainty.
Next, we obtain planet parameters for the exoplanets in systems for which an isochrones stellar property inference was successful. We select our sample of exoplanet system measurements from the NASA Exoplanet Archive’s ps table (Akeson et al., 2013; Christiansen et al., 2025; NASA Exoplanet Archive, 2025b). We eliminate systems discovered by pulsar timing, microlensing, and direct imaging as planet radii cannot be directly inferred for these systems. Our process for calculating planet parameters prefers to use as few Exoplanet Archive entries as possible for each planet’s orbit solution (Doppler semiamplitude $K$ and eccentricity $e$) and transit solution (orbital period $P_{\text{orb}}$, inclination $i$ or impact parameter $b$, and planet-to-star radius ratio $q$ or transit depth $\delta$). We prefer transit solutions with inclinations over impact parameters and radius ratios over transit depths as inclinations and radius ratios do not require additional transformations to obtain a mass or radius inference. In cases where a transit depth is reported instead of a planet-to-star radius ratio, we transform the transit depth radius ratio by taking its square root and transform its uncertainty in quadrature. We choose the larger of the upper and lower uncertainty for each parameter in cases where they are unequal as the inputs to our hierarchical Bayesian inference must be normally distributed.
@@
In a similar fashion to Schmidt et al. (2023), we re-parameterize Equation (6) as a cubic equation
@@
IV Hierarchical Bayesian Inference
@@
$^{i}, 𝐡\mathbf{h}$ Heat factor (see Section IV)
@@
Note. — The superscript $i$ indicates the $i^{\mathrm{th}}$ planet in our sample, and bolded variables are the vector of that parameter for all planets. The covariance $\mathbf{\Sigma_{*}^{i}}$ is output by the isochrones run for each host star. Additionally, as the MIST grids used by isochrones lack the initial bulk metallicity column, we fit for this using the stellar age, mass, and present-day photospheric metallicity to obtain the initial bulk metallicity for each system (see Section IV.1 for more details).
To handle possible variations in the heating efficiency $\epsilon(F)$ (Eq. 3), we parameterize the variation with $h^{i}$, which is the offset relative to the mean heating in number of standard deviations. To match the results of Thorngren and Fortney (2018) exactly, we could set the prior as $h\sim\mathcal{N}(0,1^{2})$ where $\mathcal{N}(\mu,\sigma^{2})$ is a normal distribution with mean $\mu$ and standard deviation $\sigma$. However, to allow the model additional flexibility while maintaining the functional form of $\epsilon(F)$ we instead allow a global mean offset to the heating $H$, such that $h^{i}\sim\mathcal{N}(H,1^{2})$. Due to some differences in equation of state between this article and Thorngren and Fortney (2018), we expect $H$ to be somewhat negative (i.e. less heating is required).
@@ IV.1 Sampling Strategy
Note. — This table is sorted by planet name and is available in its entirety in machine-readable format. The sample enumerated by this table has not been vetted for population-level biases and should not be used for occurrence analyses.
Our thermal evolution model’s stellar model grid is adapted from the MIST grid and takes the star’s initial bulk metallicity as an input. However, the MIST model grid used by the isochrones Python package is restricted to a subset of parameters that does not include it; as a result the posteriors instead report the present-day photospheric metallicity. This requires us to use a root finder to calculate the initial bulk metallicity posterior for each star from its inferred present-day photospheric metallicity, stellar mass, and stellar age posteriors. While this process works with no issue for the vast majority of the host stars in our sample, some metal-rich stars lack suitable MIST models at a high enough initial bulk metallicity to match the inferred present-day photospheric metallicity from their isochrones posterior. Performing the initial bulk metallicity inference procedure cuts off a significant portion of these stars’ metallicity posteriors. We therefore exclude 33 stars for which this occurs to more than a few percent of samples.
We perform individual interior structure retrievals on all of the planets in our sample for which our model is suitable: 0.1 $_{\text{Jup}}<M^{i}_{\text{obs}}<4~M$$_{\text{Jup}}$, 0.5 $_{\odot}<M^{i}_{\ast\text{obs}}<2~M$, and $0.75<\text{[Fe/H]}^{i}_{\text{0}}<0.5$, totaling 273. Though we compute our model for each iteration, underlying it is a static model grid with the following parameter ranges and resolutions:
  1. 1.
    $M^{i}$, with 114 steps between 1.0050608060184631 and 9534.844001949032 M;
  2. 2.
    $Z^{i}$, with 21 steps between 0 and 0.9999;
  3. 3.
    $f_{\text{core}}$, the fraction of mass in the core, with 21 steps between 0 and 1; and
  4. 4.
    $s_{i}$, with 225 steps between 5 and 14.
These resulting posterior distributions are often non-Gaussian, so to facilitate convergence we utilize a two-step MCMC process. We first run a shorter 20,000-iteration initial retrieval with a 5000-iteration burn-in to characterize the covariance, and then run a second 250,000-iteration retrieval with a 5000-iteration burn-in and using the covariance of the initial run to set the proposal covariance. We apply a thinning factor of 10 to both of these steps, where we only record a sample every 10 iterations. In some cases the model fails to converge or otherwise fails to provide plausible results; we exclude these planets from the upper level of the hierarchical model. We list the planets in our sample for which we successfully retrieve the interior structure following this methodology in Table 2.
@@
V Results and Discussion
We show an example corner plot from retrievals on WASP-6 b (green) and CoRoT-25 b (blue) in Figure 3. Our model allows a relatively young planet such as WASP-6 b to be larger due to either delayed cooling or a less dense envelope caused by a lower bulk metallicity. This manifests in the posterior as a degeneracy between the planet’s bulk metallicity $Z_{p}$ and intrinsic luminosity reduction factor $f_{\text{L}}$. On the other hand, a planet like CoRoT-25 b (Fig. 3, right) is too old to be affected by delayed cooling, having already reached thermal equilibrium where its size is coupled to its host star’s evolution (see Sec. II). Such planets are immune to the $Z_{p}$-$f_{\text{L}}$ degeneracy, making them a natural control group for delayed cooling and useful for constraining the population mean heat factor $H$. These examples show that a single planet cannot provide enough information to fully constrain $f_{\text{L}}$, making a population-level analysis necessary.
@@ V Results and Discussion
Next, we show the results of our hierarchical Bayesian inferences in Figure 4. Each corner plot in Figure 4 shows the posterior and covariance for our population mean heat factor $H$ and intrinsic luminosity reduction factor $f_{\text{L}}$. These parameters are uncorrelated, indicating that the delayed cooling effect we see is not biased by or an artifact of our thermal evolution model’s equation of state choice. We vary the value of each retrieved planet’s initial specific entropy from $s_{\text{init}}=11~k_{B}/\text{baryon}$ to $s_{\text{init}}=9~k_{B}/\text{baryon}$ in increments of $0.5~k_{B}/\text{baryon}$. These represent the range between “hot start” and “cold start” giant planet formation models (Marley et al., 2007) to avoid limiting our modeling to a particular assumption about how the planet initially formed. We note that the shape of the posterior for the $s_{\text{init}}=11~k_{B}/\text{baryon}$ case is more jagged in comparison the other posteriors; we attribute this to the greater influence of the youngest planets in our sample on it, with their samples being preferred more strongly compared to the posteriors with the lower values of $s_{\text{init}}$. Regardless, our results are consistent with one another for each value of $s_{\text{init}}$.
@@
Figure 4: Corner plots for the hyperparameters of our population-level intrinsic luminosity reduction factor $f_{\text{L}}$ and population mean heat factor $H$. Each posterior’s model assumes a different initial specific entropy for the individual interior structure retrievals it marginalizes over. Not shown are the individual planets’ heat factor hyperparameters, which are included to account for small differences in the theoretical equation of state used between this work and Thorngren and Fortney (2018). We find that the interior cooling rates for hot Jupiters are reduced by a factor of 20–60 relative to Thorngren and Fortney (2018), regardless of our assumption for the initial specific entropy.
@@
We find that the interior cooling rates of the hot Jupiter population are reduced by a factor of between 20 and 60 relative to the simpler heating model presented in Thorngren and Fortney (2018). In cases where the planet has reached thermal equilibrium, this is equivalent to shallow heating accounting for about 97% of the total heating experienced by hot Jupiters [^6] Note that while most of the planets in our sample are not exactly in equilibrium, the older ones are not far off from it.. In other words, the delayed cooling we observe in the hot Jupiter population means that the youngest ($t^{i}<$1 Gyr) planets in our sample are consistently larger than the trend inferred by Thorngren and Fortney (2018). We additionally find that the population mean heat $H$ (see Sec. IV) is modest but nonzero to correct for the updated equation of state. This means that the base amount of heating required to inflate hot Jupiters’ radii is slightly lower than Thorngren and Fortney (2018) [^7] Also note that, due to the 3D nature of hot Jupiters, the amount of heating we infer here is likely an underestimate (Zhang, 2023).. The lack of correlation between $H$ and $f_{\text{L}}$ in our posteriors (Fig. 4) shows that this was not a bias on the latter.
@@
The high intrinsic luminosity reduction factor we find is best explained by predominantly shallow heating (near the surface, rather than the deep interior). The proposed hot Jupiter heating mechanisms that would cause shallow heating, including Ohmic dissipation (Batygin and Stevenson, 2010; Batygin et al., 2011), turbulent mixing (Youdin and Mitchell, 2010), and longitudinal advection (Tremblin et al., 2017) preferentially occur near the radiative–convective boundary (RCB). The exact location of this heating with respect to the RCB is only of minor importance; while shallow heating mechanisms that operate above it can delay cooling by shifting it deeper into the interior (e.g., Youdin and Mitchell, 2010), very shallow heating cannot cause reinflation on main sequence or post-main sequence timescales (Komacek et al., 2020) and is unable to significantly delay cooling (Spiegel and Burrows, 2013). While a larger-than-expected opacity in the shallow interior (Burrows et al., 2003, 2007) may also contribute, the extent to which our hot Jupiters’ radii are inflated at the population level indicates that this explanation alone is insufficient. Similarly, strong compositional gradients (e.g., Chabrier and Baraffe, 2007; Leconte and Chabrier, 2012) would be insufficient, as neither they nor enhanced opacities could explain the flux-heating relation of Thorngren and Fortney (2018). As a result, while we cannot definitively rule out these alternatives entirely, they can only be minor contributions compared to shallow heating mechanisms.
@@
Our results strongly exclude $f_{\text{L}}=1$, meaning some delay is required to explain the larger radii of young hot Jupiters. On the other hand, exclusively delayed cooling effects from shallow heating would not reinflate planets as seen in hot Jupiters orbiting main sequence (Hartman et al., 2016) and evolved (Grunblatt et al., 2016, 2017) stars. Our model’s preferred values of $f_{\text{L}}$, a few percent of standard cooling rates, is able to reinflate hot Jupiters with over an order of magnitude less deep heating power than the amount predicted by simpler hot Jupiter anomalous heating models. This balances the need for both inflated young hot Jupiters and reinflation of older ones, resolving the tension between the deep heating models of Thorngren and Fortney (2018) and the delayed cooling seen in Thorngren et al. (2021). It may also ease the tension between deep heating models and the observed population (e.g., Leconte et al., 2010; Yee and Vissapragada, 2025, in the case of tidal dissipation), as much less heating of this kind is necessary to produce the same effect.
@@
Dominant shallow heating has implications for the relationship between circulation-related exoplanet observables and radius inflation. This is because the strength of many shallow heating mechanisms is coupled to wind speeds in the atmosphere (see Menou, 2012), so circulation-related observables should mirror the heating efficiency-equilibrium temperature trend found in Thorngren and Fortney (2018). Specifically, we expect hot spot offsets for hot Jupiters to scale with the log of the incident flux as $\exp{\left[(\log(F)-\log(F_{0}))^{2}/(2w)\right]}$, where $F_{0}\approx 1.38$ Gerg s-1 cm-2 and $w\approx 0.37$ (Thorngren, 2024). This corresponds to a peak hot spot offset at $T_{\mathrm{eq}}\approx 1600$ K, decreasing again at higher temperatures. However, as phase curve offsets are also impacted by the propagation of planetary-scale Kelvin and Rossby waves independent of wind speeds (Lewis and Hammond, 2022), confirmation of this relationship will require a large body of high-quality phase curve data to constrain wind speeds. Though some population studies of archival Spitzer phase curves exist (e.g., May et al., 2022; Dang et al., 2025), they are limited in sample size as well as the range of equilibrium temperatures and instellations they probe. Future observing campaigns, as well as missions like European Space Agency’s Ariel mission (Tinetti et al., 2022), could observe a larger statistical sample to evaluate these predictions. Other circulation-related observables, such as bulk wind speeds inferred from high-precision Doppler spectroscopy of hot Jupiter atmospheres (Seidel et al., 2025), should also exhibit a similar trend as found by Thorngren and Fortney (2018).
@@
While our sample was large enough to obtain definitive results, a larger sample of younger ($t^{i}_{\text{obs}}\lesssim 1$ Gyr) planets with precise masses, radii, and host star properties would enable testing of more complex thermal evolution models. As hot Jupiter systems experience a range of formation and evolutionary histories, the value of $f_{\text{L}}$ for a given hot Jupiter may not always be constant or equal for both the heating and cooling terms. The careful study of higher-order effects due to tidal circularization and inspiral (e.g., Jackson et al., 2008; Leconte et al., 2010) and/or late-arrival scenarios (e.g., Hamer and Schlaufman, 2022; Schmidt and Schlaufman, 2026) further altering the planet’s interior heating or instellation could provide more physically-motivated constraints on delayed cooling. In the particular case of late arrivals, though the evolution of a hot Jupiter’s orbit in the era of its parent protoplanetary disk can impact its subsequent thermal and radius evolution (Mol Lous and Miguel, 2020), high-eccentricity tidal migration can effectively “reset” the thermal evolution of a late-arriving hot Jupiter that cooled and contracted far from its host star for over a Gyr (e.g., Liu et al., 2008; Ibgui and Burrows, 2009). On the other end of the exoplanet age distribution, a larger sample of hot Jupiters orbiting evolved stars could also help constrain the timescale for reinflation and therefore the effects of deep heating thanks to their much faster rate of stellar evolution. These larger and more precise data sets would enable further characterization of the mechanisms that heat hot Jupiters, leading to much more informed constraints on individual mechanism contributions as well as their typical impacts on observed planetary atmosphere characteristics.
VI Conclusions